# Multi-Source Position and Direction-of-Arrival Estimation Based on Euclidean Distance Matrices

### v1

@@ V Baseline Localization Methods
In this section, we briefly review SRP-based methods to localize multiple sources, both for position estimation (Section V-A) and DOA estimation (Section V-B). These methods will be used as baseline methods for the experimental evaluation in the next section. Contrary to the EDM-based methods proposed in Sections III and IV, SRP-based methods require joint optimization of a functional over multiple continuous variables.
@@ V-A SRP-Based Multi-Source Position Estimation
Similarly to the GCC-PHAT function in (20), the SRP-PHAT functional for the 3-dimensional position estimation [22] is defined as
@@
where $\tau_{ij}(\mathbf{p})$ denotes the TDOA (8) corresponding to position $\mathbf{p}=[p_{x},p_{y},p_{z}]^{\textrm{T}}$, and the summation considers all microphone pairs, where $i>j$. The estimated source positions $\hat{\mathbf{p}}_{1},\dots,\hat{\mathbf{p}}_{S}$ are determined as the vectors that correspond to the $S$ largest local maxima of the SRP-PHAT functional. This requires a joint optimization of three continuous variables for all feasible positions within the room boundaties ($P_{x}^{\textrm{min}}\leq p_{x}\leq P_{x}^{\textrm{max}}$, $P_{y}^{\textrm{min}}\leq p_{y}\leq P_{y}^{\textrm{max}}$, and $P_{z}^{\textrm{min}}\leq p_{z}\leq P_{z}^{\textrm{max}}$). Practical considerations regarding the discretization of the continuous position variables and the determination of local maxima are discussed in Section VI.
@@ V-B SRP-Based Multi-Source DOA Estimation
Similarly to (57), the SRP-PHAT functional for DOA estimation is defined as
@@
where $\tau_{ij}(\mathbf{v})=(\mathbf{m}_{i}-\mathbf{m}_{j})^{\text{T}}\mathbf{v}/\nu$ denotes the TDOA corresponding to the DOA vector $\mathbf{v}$, which depends on the azimuth and elevation angles. The estimated azimuth and elevation angles $\hat{\theta}_{1},\dots,\hat{\theta}_{S}$ and $\hat{\phi}_{1},\dots,\hat{\phi}_{S}$ are determined from the DOA vectors that correspond to the $S$ largest local maxima of the SRP-PHAT functional. This requires a joint optimization of two continuous variables for all feasible azimuth angles $\pi<\theta\leq\pi$ and elevation angles $\pi/2<\phi\leq\pi/2$. Practical considerations regarding the discretization of the azimuth and elevation angles and the determination of local maxima are discussed in Section VI.
@@ VI Experimental Evaluation
In this section, we compare the performance of the proposed EDM-based methods with the baseline SRP-based methods for multi-source position and DOA estimation in noisy and reverberant environments. We conducted two sets of experiments, where in experiment 1 we evaluated 3D position estimation using spatially distributed microphones and in experiment 2 we evaluated 3D DOA estimation using compact microphone arrays. Section VI-A describes the acoustic scenarios for both experiments. The implementation details of the EDM-based and SRP-based methods are presented in Sections VI-B and VI-C, respectively. The results of experiment 1 (position estimation) are discussed in Section VI-D, while the results of experiment 2 (DOA estimation) are discussed in Section VI-E.
@@ VI-A Acoustic Scenarios
In experiment 1, the spatially distributed microphones were positioned randomly for each scenario within a cube with sides of length 2 m, with a minimum distance of 10 cm between the microphones. In experiment 2, the microphones were positioned randomly for each scenario within a smaller cube with sides of length 10 cm, with a minimum distance of 4 cm between microphones. In both experiments, the distance between the second source and the centroid of the microphone array was fixed at $d_{\textrm{c}2}=2$ m. To evaluate the influence of source distance on localization accuracy, different distances $d_{\textrm{c}1}$ between the first source and the centroid of the microphone array were considered. In experiment 1, we considered $d_{\textrm{c}1}\in\{0,1,2,3,4\}$ m, where $d_{\textrm{c}1}=0$ m and $d_{\textrm{c}1}=1$ m correspond to the first source being inside of the cube, while $d_{\textrm{c}1}=3$ m and $d_{\textrm{c}1}=4$ m correspond to the first source being outside of the cube. In experiment 2, we considered $d_{\textrm{c}1}\in\{0.5,1,2,3,4\}$ m, corresponding to both sources being outside of the cube (in the far field of the microphone array). For all scenarios, both sources were spaced at least 1 m apart and maintained a minimal angular separation of $20^{\circ}$ relative to the array centroid.
@@ VI-B Implementation of EDM-Based Source Localization
As explained in Sections III and IV, the proposed EDM-based localization methods require candidate TDOA estimates, which are obtained from the GCC-PHAT function. In practice, the time-domain microphone signals are first transformed to the short-time Fourier transform (STFT) domain, with a frame length of $K=512$ samples (corresponding to 32 ms), $50\%$ overlap between frames, and using a square-root-Hann analysis window. Similarly to (21), the instantaneous normalized phase spectrum between the $i$-th and $j$-th microphones is computed as
@@
where $Y_{i}[k,l]$ denotes the STFT coefficient of the $i$-th microphone signal at frequency bin $k\in\{0,\dots,K-1\}$ and time frame $l\in\{1,\dots,L\}$, where $L$ denotes the number of frames. Similarly to (20), the discrete-time GCC-PHAT function between the $m$-th microphone and the reference microphone is computed as
@@
where $n$ denotes the discrete-time index, with $\tau=n/f_{s}$. To achieve a more precise TDOA estimate, the discrete-time GCC-PHAT function $\xi_{m}[n,l]$ is interpolated by a factor $R=20$ using resampling. Only plausible time-lags $n_{m}$ between the $m$-th microphone and the reference microphone are considered, determined by the inter-microphone distance $D_{m1}$, i.e., $|n_{m}|<Rf_{s}D_{m1}/\nu$. To emphasize strong peaks, the function is weighted as $\tilde{\xi}_{m}[n_{m},l]=\exp{(\gamma\xi_{m}[n_{m},l])}$, with $\gamma=30$ for position estimation (experiment 1) and $\gamma=50$ for DOA estimation (experiment 2). Since the sources are assumed to be static, the GCC-PHAT function is averaged over all frames, i.e.,
@@
For EDM-based position estimation, the optimal distance variable in (28) was first determined using a one-dimensional grid search, with a resolution of 1 cm and a maximum distance of 6 m (i.e., 601 grid points) and then fine-tuned with a quadratic interpolation. $C=3$ candidate TDOA estimates are considered per microphone pair, resulting in $Q=C^{M-1}=243$ total combinations. After computing the EDM-based cost functions $J(\alpha,q)$ in (26) and determining the optimal distance variable $\hat{\alpha}(q)$ using a one-dimensional grid search for all $Q$ possible combinations of candidate TDOA estimates, the respective cost function minima were sorted. The combination $\hat{q}_{1}$ yielding the smallest cost function minimum $J(\hat{\alpha}(\hat{q}_{1}),\hat{q}_{1})$ was used to estimate the source position $\hat{\mathbf{p}}_{1}$. To estimate the second source position $\hat{\mathbf{p}}_{2}$, only combinations where at least four candidate TDOA estimates differ were considered, as the likelihood that more than three TDOAs are shared between sources was assumed to be negligible.
For EDM-based DOA estimation, only $C=2$ candidate TDOA estimates were considered, resulting in $Q=32$ total combinations. Unlike with spatially distributed microphones, preliminary experiments with compact microphone arrays indicated that choosing $C>2$ did not notably improve the accuracy of the localization accuracy, since the number of spurious peaks in the GCC-PHAT function is typically quite low when considering plausible time-lags for compact microphone arrays. After computing the EDM-based cost function values $I(q)$ in (53) for all $Q$ possible combinations of candidate TDOA estimates, the respective cost function values were sorted. The combination $\hat{q}_{1}$ yielding the smallest value $I(\hat{q}_{1})$ was used to estimate the source DOA $\hat{\mathbf{v}}_{1}$. To estimate the second source DOA $\hat{\mathbf{v}}_{2}$, only combinations where at least four candidate TDOA estimates differ were considered.
It is known that the choice of reference microphone affects the TDOA estimation accuracy for single source scenarios [44]. Since preliminary experiments using multiple sources also demonstrated that the localization accuracy of the EDM-based methods may be negatively influenced by a poor choice of reference microphone, the reference microphone was chosen depending on the array geometry. For spatially distributed microphone arrays (experiment 1), the reference microphone was chosen as the microphone closest to the array centroid. This reduces the likelihood that the reference microphone is located at a large distance from one or both sources, which would result in a poor SNR. For compact microphone arrays (experiment 2), the reference microphone was chosen as the microphone which was on average farthest from the other microphones, aiming at mitigating the influence of reverberation and noise on the reliability of the GCC-PHAT function [45].
@@ VI-C Implementation of SRP-Based Source Localization
For the SRP-based localization methods (Section V), the exhaustive search over continuous variables is implemented as a discretized grid search. Using the instantaneous normalized phase spectrum in (59), the SRP-PHAT functionals for position and DOA estimation in (57) and (58), are computed as
@@
Similarly as for the GCC-PHAT function in (61), the SRP-PHAT functionals are averaged over all frames, i.e.,
$\Psi({\mathbf{p}})\;=\;\frac{1}{L}\sum_{l=1}^{L}\Psi[l]({\mathbf{p}})\;\;,$ (64)
$\Psi({\mathbf{v}})\;=\;\frac{1}{L}\sum_{l=1}^{L}\Psi[l]({\mathbf{v}})\;\;.$ (65)
@@
Exhaustively searching for the local maxima of these functionals can be computationally demanding if the grid resolution is high (especially for three-dimensional position estimation). Therefore, we first compute $\beta\geq S$ candidate positions/DOAs on a coarse grid (similarly to [46]) that spans the entire range of feasible positions/DOAs. We then refine these coarse estimates by re-evaluating the functional on a finer grid in the vicinity of the coarse estimates. Finally, the $S$ positions $\hat{\mathbf{p}}_{1}^{\textrm{SRP}},\dots,\hat{\mathbf{p}}_{S}^{\textrm{SRP}}$ or DOAs $\hat{\mathbf{v}}_{1}^{\textrm{SRP}},\dots,\hat{\mathbf{v}}_{S}^{\textrm{SRP}}$ are estimated as those corresponding to the $S$ highest SRP-PHAT values on the finer grids.
@@
For position estimation, the functional in (64) was first evaluated on a coarse grid with a 10 cm resolution along the x-, y-, and z-axes (i.e., $59\times 59\times 23=80,063$ grid points). The $\beta=3$ coarse grid points with the highest SRP-PHAT values were then re-evaluated on a finer grid with a 1 cm resolution around those points, within a 20 cm $\times$ 20 cm $\times$ 20 cm cube (i.e., $3\times 21^{3}=27,783$ grid points). After estimating the first source position $\hat{\mathbf{p}}_{1}^{\textrm{SRP}}$ corresponding to the largest value of the fine grids, for the second source only coarse grid points at least 50 cm from the coarse estimate corresponding to the first source position were considered to avoid considering high SRP-PHAT values in the neighbourhood of $\hat{\mathbf{p}}_{1}^{\textrm{SRP}}$ as separate sources (which could sometimes result in additional grid points being considered). For DOA estimation, the functional in (65) was first evaluated on a coarse grid with a $5^{\circ}$ resolution in azimuth ($\theta$) and elevation ($\psi$) (i.e., $(72\times 35)+2=2,522$ grid points - noting that for elevation angles 0 and 180 the DOA is independent of the azimuth angle). The $\beta=2$ coarse grid points with the highest SRP-PHAT values were then re-evaluated on a finer grid with a $0.5^{\circ}$ resolution, within $10^{\circ}$ in azimuth and elevation angles (i.e., $2\times 21^{2}=882$ grid points). After estimating the azimuth $\hat{\theta}_{1}^{\textrm{SRP}}$ and elevation $\hat{\psi}_{1}^{\textrm{SRP}}$ of the first source, for the second source only coarse grid points at least $20^{\circ}$ away from the coarse estimate corresponding to the first source DOA were considered (sometimes resulting in an additional coarse grid point being considered).
VI-D Experiment 1: Position Estimation
This section compares the position estimation accuracy of the proposed EDM-based method and the baseline SRP-based method for the acoustic scenarios with spatially distributed microphones described in Section VI-A. To evaluate the performance, we consider the position estimation error for each source, defined as
[…]
where the estimated source positions $\hat{\mathbf{p}}_{1}$ or $\hat{\mathbf{p}}_{2}$ are assigned to true source positions ${\mathbf{p}}_{1}$ or ${\mathbf{p}}_{2}$ using greedy assignment [47] (i.e., assigning each estimated source to the closest true source sequentially, in order of increasing $\varepsilon_{s}^{\textrm{pos}}$). Fig. 5 presents box plots of the position estimation errors $\varepsilon_{1}^{\textrm{pos}}$ and $\varepsilon_{2}^{\textrm{pos}}$ over 100 simulated scenarios, considering different distances $d_{\textrm{c}1}$ between source 1 and the array centroid.
Figure 5: Box plots of position estimation errors for both sources. The position estimation errors are shown for the EDM-based and SRP-based methods, for different distances $d_{\textrm{c}1}$ between source 1 and the array centroid ($d_{\textrm{c}2}=2$ m, 2 m cube). The red numbers at the top denote the number of results outside of the plotted range.
Table I summarizes the median position estimation errors for both sources.
TABLE I: Median position estimation errors for the EDM-based and SRP-based methods for different distances $d_{\textrm{c}1}$ between source 1 and the array centroid.
Median position estimation error [cm]
$d_{\textrm{c}1}[m]$ EDM, $\overline{\varepsilon}_{1}^{\textrm{pos}}$ EDM, $\overline{\varepsilon}_{2}^{\textrm{pos}}$ SRP, $\overline{\varepsilon}_{1}^{\textrm{pos}}$ SRP, $\overline{\varepsilon}_{2}^{\textrm{pos}}$
0 0.1 0.8 58.0 197.2
11 0.2 0.7 4.1 136.1
22 0.6 0.7 6.4 7.7
33 2.3 0.6 200.0 3.7
44 9.4 0.6 350.6 3.1
As can be observed from Table I, for all source distances $d_{\textrm{c}1}$ the proposed EDM-based method yields considerably lower median position estimation errors for both sources than the SRP-based method. In addition, it can be observed from Fig. 5 that the spread of the box plots is much smaller for the EDM-based method than for the SRP-based method for all source configurations. Furthermore, it is interesting to note that both methods behave quite differently for different source configurations. As can be observed from Table I, the median position estimation error for the EDM-based method mainly depends on source distance, i.e., increases for source 1 with increasing $d_{\textrm{c}1}$, whereas it remains relatively constant for source 2 (at $d_{\textrm{c}2}=2$ m). On the other hand, the median position estimation error for the SRP-based method depends both on the absolute distances between the sources and the array centroid as well as on the relative distance between the sources. For instance, it can be clearly observed that the median position estimation error is larger for the source that is farther away from the array centroid (e.g., for $d_{\textrm{c}1}=3$ m, $\overline{\varepsilon}_{1}^{\textrm{pos}}(\mathbf{s})=200$ cm and $\overline{\varepsilon}_{2}^{\textrm{pos}}(\mathbf{s})=3.7$ cm). This can be explained by the peaks of the farther source in the SRP functional being overpowered by the peaks of the closer source. In addition, when both sources are close to the microphones, the median position estimation error for both sources is large (e.g., for $d_{\textrm{c}1}=0$ m, $\overline{\varepsilon}_{1}^{\textrm{pos}}(\mathbf{s})=58$ cm and $\overline{\varepsilon}_{2}^{\textrm{pos}}(\mathbf{s})=197.2$ cm). This corresponds to [21, 24], where it was shown that the accuracy of SRP-based position estimation degrades for sources located close to the microphone array because the peaks in the SRP-PHAT functional become narrow. This would necessitate a high 3D grid resolution to detect these peaks, which is computationally impractical in most applications. In contrast, the EDM-based method only requires a one-dimensional grid search on the distance variable $\alpha$, which is computationally more efficient to optimize at a high resolution.
The average run times of the EDM-based and SRP-based position estimation methods are shown in Table II, when processing 5 second signals on a Ryzen 5900x processor. The EDM-based method is about 380 times faster than the SRP-based method, which can be attributed to the vast difference in search-space between both methods. Rather than jointly optimizing a functional in three position variables using all microphone pairs, the EDM-based method optimizes a cost function in a single continuous variable, based on estimated TDOAs between the reference microphone and the other microphones.

### v2

@@ V Baseline Localization Methods
In this section, we briefly review SRP-based methods to localize multiple sources, both for position estimation (Section V-A) and DOA estimation (Section V-B). These methods will be used as baselines in the experimental evaluation in the next section. Contrary to the EDM-based methods proposed in Sections III and IV, SRP-based methods require joint optimization of a functional over multiple continuous variables.
@@ V-A SRP-Based Multi-Source Position Estimation
Similarly to the GCC-PHAT function in (16), the SRP-PHAT functional for 3-dimensional position estimation [23] is defined as
@@
where $\tau_{ij}(\mathbf{p})$ denotes the TDOA (4) corresponding to position $\mathbf{p}=[p_{x},p_{y},p_{z}]^{\textrm{T}}$, and the summation considers all microphone pairs, where $i>j$. The estimated source positions $\hat{\mathbf{p}}_{1},\dots,\hat{\mathbf{p}}_{S}$ are determined as the vectors that correspond to the $S$ largest local maxima of the SRP-PHAT functional. This requires a joint optimization of three continuous variables for all feasible positions within the room boundaties ($P_{x}^{\textrm{min}}\leq p_{x}\leq P_{x}^{\textrm{max}}$, $P_{y}^{\textrm{min}}\leq p_{y}\leq P_{y}^{\textrm{max}}$, and $P_{z}^{\textrm{min}}\leq p_{z}\leq P_{z}^{\textrm{max}}$). Practical considerations regarding the discretization of the continuous position variables and the determination of local maxima are discussed in Section VI-C.
@@ V-B SRP-Based Multi-Source DOA Estimation
Similarly to the SRP-PHAT function for position estimation, the SRP-PHAT functional for DOA estimation is defined as
@@
where $\tau_{ij}(\mathbf{v})=(\mathbf{m}_{i}-\mathbf{m}_{j})^{\text{T}}\mathbf{v}/\nu$ denotes the TDOA corresponding to the DOA vector $\mathbf{v}$, which depends on the azimuth and elevation angles. The estimated azimuth and elevation angles $\hat{\theta}_{1},\dots,\hat{\theta}_{S}$ and $\hat{\phi}_{1},\dots,\hat{\phi}_{S}$ are determined from the DOA vectors that correspond to the $S$ largest local maxima of the SRP-PHAT functional. This requires a joint optimization of two continuous variables for all feasible azimuth angles $\pi<\theta\leq\pi$ and elevation angles $\pi/2<\phi\leq\pi/2$. Practical considerations regarding the discretization of the azimuth and elevation angles and the determination of local maxima are discussed in Section VI-C.
@@ VI Experimental Evaluation
In this section, we compare the performance of the proposed EDM-based methods with the baseline SRP-based methods for multi-source position and DOA estimation in noisy and reverberant environments. We conducted two sets of experiments, where in experiment 1 we evaluated 3D position estimation using spatially distributed microphones and in experiment 2 we evaluated 3D DOA estimation using compact microphone arrays. Section VI-A describes the acoustic scenarios for both experiments. The implementation details of the EDM-based and SRP-based methods are presented in Sections VI-B and VI-C, respectively, and their computational complexity is compared in Section VI-D. The results of experiment 1 (position estimation) are discussed in Section VI-E, while the results of experiment 2 (DOA estimation) are discussed in Section VI-F.
@@ VI-A Acoustic Scenarios
In experiment 1, the spatially distributed microphones were positioned randomly for each scenario within a cube with sides of length 2 m, with a minimum distance of 10 cm between microphones. In experiment 2, the microphones were positioned randomly for each scenario within a smaller cube with sides of length 10 cm, with a minimum distance of 4 cm between microphones. In both experiments, the distance between the second source and the centroid of the microphone array was fixed at $d_{\textrm{c}2}=2$ m. To evaluate the influence of source distance on localization accuracy, different distances $d_{\textrm{c}1}$ between the first source and the centroid of the microphone array were considered. In experiment 1, we considered $d_{\textrm{c}1}\in\{0,1,2,3,4\}$ m, where $d_{\textrm{c}1}=0$ m and $d_{\textrm{c}1}=1$ m correspond to the first source being inside of the cube, while $d_{\textrm{c}1}=3$ m and $d_{\textrm{c}1}=4$ m correspond to the first source being outside of the cube. In experiment 2, we considered $d_{\textrm{c}1}\in\{0.5,1,2,3,4\}$ m, corresponding to both sources being outside of the cube (in the far field of the microphone array). For all scenarios, both sources were spaced at least 1 m apart and maintained a minimal angular separation of $20^{\circ}$ relative to the array centroid. Exemplary simulated scenarios for both experiments are plotted in Fig. 7 (supplementary material).
@@ VI-B Implementation of EDM-Based Source Localization
As explained in Sections III and IV, the proposed EDM-based localization methods require candidate TDOA estimates, which are obtained from the GCC-PHAT function. In practice, the time-domain microphone signals are first transformed to the short-time Fourier transform (STFT) domain, with a frame length of $K=512$ samples (i.e., 32 ms), $50\%$ overlap between frames, and using a square-root-Hann analysis window. Similarly to (17), the instantaneous normalized phase spectrum between the $i$-th and $j$-th microphones is computed as
@@
where $Y_{i}[k,l]$ denotes the STFT coefficient of the $i$-th microphone signal at frequency bin $k\in\{0,\dots,K-1\}$ and time frame $l\in\{1,\dots,L\}$, where $L$ denotes the number of frames (i.e., $L=313$ frames for a 5 second signal). Similarly to (16), the discrete-time GCC-PHAT function between the $m$-th and the reference microphone is computed as
@@
where $n$ denotes the discrete-time index, with $\tau=n/f_{s}$. To achieve a more precise TDOA estimate, the discrete-time GCC-PHAT function $\xi_{m}[n,l]$ is interpolated by a factor $R=20$ using resampling. Only plausible time-lags $n_{m}$ between the $m$-th and the reference microphone are considered, determined by the inter-microphone distance $D_{m1}$, i.e., $|n_{m}|<Rf_{s}D_{m1}/\nu$. To emphasize strong peaks, the function is weighted as $\tilde{\xi}_{m}[n_{m},l]=\exp{(\gamma\xi_{m}[n_{m},l])}$, with $\gamma=30$ for position estimation (experiment 1) and $\gamma=50$ for DOA estimation (experiment 2). Since the sources are assumed to be static, the GCC-PHAT function is averaged over all frames, i.e.,
@@
For EDM-based position estimation, the optimal distance variable in (24) was first determined using a one-dimensional grid search, with a resolution of 1 cm and a maximum distance of 6 m (i.e., $G_{\alpha}=601$ grid points) and then fine-tuned with a quadratic interpolation. $C\in\{2,3\}$ candidate TDOA estimates were considered per microphone pair, resulting in $Q=C^{M-1}\in\{32,243\}$ total combinations, respectively. After computing the EDM-based cost functions $J(\alpha,q)$ in (22) and determining the optimal distance variable $\hat{\alpha}(q)$ using a one-dimensional grid search for all $Q$ possible combinations of candidate TDOA estimates, the respective cost function minima were sorted. The combination $\hat{q}_{1}$ yielding the smallest cost function minimum $J(\hat{\alpha}(\hat{q}_{1}),\hat{q}_{1})$ was used to estimate the source position $\hat{\mathbf{p}}_{1}$. To estimate the second source position $\hat{\mathbf{p}}_{2}$, only combinations where at least four candidate TDOA estimates differ were considered, as the likelihood that more than three TDOAs are shared between sources was assumed to be negligible. This rule prevents the EDM-based method from detecting sources that are spatially close.
For EDM-based DOA estimation, $C\in\{2,3\}$ candidate TDOA estimates were considered per microphone pair, resulting in $Q\in\{32,243\}$ total combinations, respectively. After computing the EDM-based cost function values $I(q)$ in (48) for all $Q$ possible combinations of candidate TDOA estimates, the respective cost function values were sorted. The combination $\hat{q}_{1}$ yielding the smallest value $I(\hat{q}_{1})$ was used to estimate the source DOA $\hat{\mathbf{v}}_{1}$. To estimate the second source DOA $\hat{\mathbf{v}}_{2}$, only combinations where at least four candidate TDOA estimates differ were considered.
It is known that the choice of reference microphone affects the TDOA estimation accuracy for single source scenarios [7]. Since preliminary experiments using multiple sources also demonstrated that the localization accuracy of the EDM-based methods may be negatively influenced by a poor choice of reference microphone, the reference microphone was chosen depending on the array geometry. For spatially distributed microphone arrays (experiment 1), the reference microphone was chosen as the microphone closest to the array centroid, reducing the likelihood that the reference microphone is located at a large distance from one or both sources which would result in a poor SNR. For compact microphone arrays (experiment 2), the reference microphone for estimating the centered TDOAs $\hat{\tau}_{m}$ in (46) was chosen as the microphone which was on average farthest from the other microphones, aiming at mitigating the influence of reverberation and noise on the reliability of the GCC-PHAT function [6].
@@ VI-C Implementation of SRP-Based Source Localization
For the SRP-based localization methods (Section V), the exhaustive search over continuous variables is implemented as a discretized grid search. Using the instantaneous normalized phase spectrum in (54), the SRP-PHAT functionals for position and DOA estimation in Sections V-A and V-B are computed as
@@
Similarly as for the GCC-PHAT function in (56), the SRP-PHAT functionals are averaged over all frames, i.e., $\Psi(\mathbf{p})=\frac{1}{L}\sum_{l=1}^{L}\Psi[l](\mathbf{p})\;,$ $\Psi(\mathbf{v})=\frac{1}{L}\sum_{l=1}^{L}\Psi[l](\mathbf{v})\;.$ (59)
@@
Exhaustively searching for the local maxima of these functionals can be computationally demanding if the grid resolution is high (especially for three-dimensional position estimation). Therefore, we first compute $\beta\geq S$ candidate positions/DOAs on a coarse grid (similarly to [14]) that spans the entire range of feasible positions/DOAs. We then refine these coarse estimates by re-evaluating the functional on a finer grid in the vicinity of the coarse estimates. Finally, the $S$ positions $\hat{\mathbf{p}}_{1}^{\textrm{SRP}},\dots,\hat{\mathbf{p}}_{S}^{\textrm{SRP}}$ or DOAs $\hat{\mathbf{v}}_{1}^{\textrm{SRP}},\dots,\hat{\mathbf{v}}_{S}^{\textrm{SRP}}$ are estimated using the $S$ highest SRP-PHAT values on the finer grids.
@@
For position estimation, the functional in (59) was first evaluated on a coarse grid with a 10 cm resolution along the x-, y-, and z-axes (i.e., $59\times 59\times 23=80,063$ grid points). The $\beta=3$ coarse grid points with the highest SRP-PHAT values were then re-evaluated on a finer grid with a 1 cm resolution around those points, within a 20 cm $\times$ 20 cm $\times$ 20 cm cube (i.e., $3\times 21^{3}=27,783$ grid points, resulting in a total of $G_{\mathbf{p}}=107,843$ grid points). After estimating the first source position $\hat{\mathbf{p}}_{1}^{\textrm{SRP}}$ corresponding to the largest value of the fine grids, for the second source only coarse grid points at least 50 cm from the coarse estimate corresponding to the first source position were considered. This rule prevents the SRP-based method from detecting sources that are spatially close. For DOA estimation, the functional in (59) was first evaluated on a coarse grid with a $5^{\circ}$ resolution in azimuth ($\theta$) and elevation ($\psi$) (i.e., $(72\times 34)+2=2,450$ grid points - noting that for elevation angles 0 and 180 the DOA is independent of the azimuth angle). The $\beta=2$ coarse grid points with the highest SRP-PHAT values were then re-evaluated on a finer grid with a $0.5^{\circ}$ resolution, within $10^{\circ}$ in azimuth and elevation angles (i.e., $2\times 21^{2}=882$ grid points, resulting in a total of $G_{\alpha}=3,332$ grid points). After estimating the azimuth $\hat{\theta}_{1}^{\textrm{SRP}}$ and elevation $\hat{\psi}_{1}^{\textrm{SRP}}$ of the first source, for the second source only coarse grid points at least $20^{\circ}$ away from the coarse estimate corresponding to the first source DOA were considered (sometimes resulting in an additional coarse grid point being considered).
VI-D Computational Complexity Analysis
This section analyzes the computational complexity of the EDM-based and SRP-based methods in terms of Big-O notation [38] and run time. For the EDM-based method, the computational complexity associated with TDOA estimation is dominated by the computation of the resampled GCC-PHAT function (55), i.e., $\mathcal{O}(KRM\log(KR))$ per frame. The complexity of the subsequent EDM-based localization steps, using the candidate TDOA estimates based on the time-averaged GCC-PHAT function in (56), are dominated by the eigenvalue decomposition with complexity $\mathcal{O}(M^{3})$, i.e., $\mathcal{O}\left(C^{M-1}M^{3}G_{\alpha}\right)$ for position estimation and $\mathcal{O}\left(C^{M-1}M^{3}\right)$ for DOA estimation. The complexity of the SRP-based methods is given by $\mathcal{O}\left(G_{\mathbf{p}}M^{2}LK\right)$ for position estimation and $\mathcal{O}\left(G_{\mathbf{v}}M^{2}LK\right)$ for DOA estimation.
It can be observed that the complexity of the EDM-based localization methods grows combinatorially with $C$ and $M$, which can become prohibitively large for a large number of microphones and/or when using many candidate TDOA estimates. In contrast, the complexity of the SRP-based methods only grow quadratically with $M$. In terms of discretized continuous variables, the complexity of the SRP-based methods depends on the resolution $G_{\mathbf{p}}$ of the three-dimensional variable $\mathbf{p}$ for position estimation and the resolution of $G_{\mathbf{v}}$ of the variable $\mathbf{v}$ for DOA estimation, which depends on the azimuth and elevation. In contrast, the complexity of the EDM-based position estimation method only depends on the resolution of the single variable $\alpha$ and the DOA estimation does not even require the discretization of a continuous variable.
For the considered experimental setup with $M=6$ microphones and using the discretized grid values $G_{\alpha},G_{\mathbf{p}}$, and $G_{\mathbf{v}}$ mentioned in the previous sections, we measured the run times of the EDM-based methods (with $C=3$) and the SRP-based methods using 5-second signals on a Ryzen 5900X processor. The average run times over all 100 simulated scenarios and 5 different distances $d_{\textrm{c}1}$ between source 1 and the array centroid are shown in Table I. For the values of the variables defined in the previous subsections, the EDM-based position estimation method is around 380 times faster than the SRP-based position estimation and the EDM-based DOA estimation method is around 23 times faster than the SRP-based method. It should be noted, that in two-dimensional scenarios the number of grid points required for SRP-based position and DOA estimation is significantly smaller, which would correspondingly reduce the complexity and run time of the SRP-based approaches.